A horse chases a pony 3 hours after the pony runs. Horse takes 4 hours to reach the pony. If the average speed of the horse is 35 kmph, what s the average speed of the pony?
Found 2 solutions by mananth, stanbon:
Answer by mananth(16949)

A horse chases a pony 3 hours after the pony runs. Horse takes 4 hours to reach the pony. If the average speed of the horse is 35 kmph, what s the average speed of the pony?
speed of pony = x kmph
distance run by pony in 3hours = 3x
speed of horse = 35 kmph
difference = 35-x
..
time taken to catch the pony = 4 hours
...
catch up time = d/r
4=3x/(35-x)
4(35-x)= 3x
140-4x=3x
7x = 140
/7
x= 20 kmph. speed of pony
...
m.ananth@hotmail.ca

Answer by stanbon(75887)

A horse chases a pony 3 hours after the pony runs. Horse takes 4 hours to reach the pony. If the average speed of the horse is 35 kmph, what s the average speed of the pony?
--------------------------
Horse DATA:
rate = 35 km/h ; time = 4 hrs : distance = rt = 35*4 = 140 km
----
Pony DATA:
distance = 140 km ; time = 7 hrs
---
Equation:
rate = distance/time = 140/7 = 20 kmh
----
Pony average rate = 20 km/h
================================
Cheers,
Stan H.